What is the y-intercept of 4(y-1)=2(^2(x+1) Note: the 2(x+1) is the exponent .
Answer by scott8148(6628)

y-intercept means x=0

4(y-1)=2^(2(0+1)) __ 4(y-1)=4 __ y-1=1 __ y=2